Given the totalcost function; TC = 3Q3 - Q2 + 150Q +50


A.           What is the fixed cost and what does rt shows?


B.           What is the variable cost function? WhatIs the variable cost at O = 2?


C.           What is the maximum output level at the minimum level of average variable cost

D.           At O = 3 what Is the trend of average cost function (is that increasing or decreasing)? Why?


1
Expert's answer
2021-10-02T10:49:08-0400
"FC=50"

"VC=3Q^3-Q^2+150Q"

"VC\vert_2=320"

"AVC=3Q^2-Q+150"

"\frac{\varDelta AVC}{\varDelta Q}=6Q-1"

"Q_{min}=\frac{1}{6}"

"\frac{\varDelta AVC}{\varDelta Q}\vert_3=17>0"

The trend of average cost function is increasing.